Suppose that market demand is given by the equation 𝑞

𝑑

=111.00−𝑝

qd=111.00−p, and market supply is given by the equation 𝑞

𝑠

=𝑝−15.00

qs=p−15.00. If the government imposes a price ceiling on this good at a price of $25.00, what would be the change in producer's surplus relative to the market equilibrium?


1
Expert's answer
2021-10-04T14:05:28-0400

"Qd=Qs\\111-p=p-15\\111+15=p+p\\126=2p\\p=63\\Q=63-15\\Q=48"


Price ceiling =25

"Qs=25-15=10\\surplus=48-10=38"